David has collected $6.50 in nickels and dimes. If she has 75 coins, how many of each coin does she have?
Answer by josmiceli(6784)

N = number of nickels
D = number of dimes
N + D = 75
N = 75 - D
650 = 5N + 10D
divide both sides by 5
130 = N + 2D
130 = 75 - D + 2D
D = 130 - 75
D = 55
N + D = 75
N = 75 - 55
N = 20
There are 20 nickels and 55 dimes
check answer
There are 75 coins-that checks
650 = 5N + 10D
650 = 5*20 + 10*55
650 = 100 + 550
650 = 650
OK